kekeman:

A county is in the shape of a rectangle that is 50 miles by 60 miles and has a population of 50,000. What is the average number of people living on each square mile of the county? Round your answer to the nearest whole number. 227 17 20 14

snowflake0531:

First you have to find the area of the whole thing, to find the unit population the area os 50*60 = 3000 3000 = 50000 1 =? simplify it to 3/50 = 1/x Cross multiply 50 = 3x x=16.6666666666666666666666 and then you can round

kekeman:

So it would be 17?

snowflake0531:

ye